### v1

@@ VI Trends in Quantum Computing Developments
TABLE IV: Company roadmaps with numbers of physical or logical qubits.
@@
Trapped-ion providers such as IonQ and Quantinuum focus on saller but higher-fidelity systems. In 2024, Quantinuum introduced a 56-qubit system that uses ion shuttling across a chip to perform operations pushing system level benchmarks such as quantum volume (QV) [cross_validating_2019, baldwin_re-examining_2022] to new record values. In 2025, the company released a 98 qubit next-generation system, which reached a point where the QV benchmark becomes too expensive in terms of classical computation to be carried out. IonQ targets 64 qubits by 2026 and 1,024 qubits by 2028 by connecting separate chips via photonic links for scalability. Their recent acquisition of smaller startup Oxford Ionics makes them the record holder in terms of fidelities with a two-qubit gate above $99.99\%$ fidelity [hughes_trapped-ion_2025] and SPAM fidelities as high as $99.9993\%$ [sotirova_high-fidelity_2024], albeit this is for an architecture involving microwave control different to their original laser-based one.
Photonic-qubit companies, notably PsiQuantum and Xanadu, pursue architectures that exploit integrated optics and telecom technologies. PsiQuantum has proposed a path to a one-million-qubit fault-tolerant computer by 2029, based on a fusion-based architecture [bartolucci_fusion-based_2023] that combines small photonic resource states in a probabilistic manner. However, major technical milestones, including the realization of large-scale, on-chip entanglement with low loss, are still outstanding. Xanadu is taking a different approach by using squeezed states of light, which are more resistant to signal loss. In 2022, the company showcased a prototype device that achieved a milestone similar to Google’s breakthrough in quantum computational advantage [b23]. In addition to hardware, Xanadu develops the open-source PennyLane platform to support hybrid quantum-classical algorithm design.
@@
For context, earlier resource estimates suggest that factoring a 2048-bit RSA modulus using a relatively direct implementation of Shor’s algorithm would require on the order of 8.4 million physical qubits [b13], assuming that they can be operated with sufficiently low error rates and long coherence times. More recent analyses, which incorporate algorithmic improvements and more efficient QEC pipelines, reduce this requirement but still call for very large, high-quality devices [b17, b27, b28].
@@
The best reported error rates (all in trapped ions) for single qubit operations are currently at the level of $10^{-7}$ [smith_single-qubit_2025], with the best two-qubit gates having reached the $10^{-5}$ range [hughes_trapped-ion_2025]. These numbers, however, are “hero values”. In many multi-qubit systems, effective error rates are much closer to $10^{-4}$ for single-qubit and $10^{-2}$ for two-qubit gates once full-system effects are included. In contrast, modern digital transistors exhibit error rates as low as $10^{-23}$ [b15], highlighting the vast gap between classical and quantum reliability.
Large-scale, fault-tolerant quantum computation therefore hinges on effective QEC. In recent years, experimental demonstrations have shown that QEC can, in principle, suppress logical error rates below those of the underlying physical qubits. For instance, a demonstration in 2025 [b16] reported an experiment in which an error metric is reduced from $10^{-2}$ to $10^{-7}$ using 72 physical qubits. However, that demonstration relied on post-processing rather than real-time feedback and did not correct all error channels. Real-time, active QEC remains experimentally demanding and has so far only been achieved in limited settings, e.g. in a trapped-ion experiment [b24].
@@ A-B Per-Run Success Probability
A-C Importance of the Quantum Signal
@@
A-D Certifying Substantial Quantum Signal in Practice
@@
A-E Shor in the Ideal Regime
@@
A-F Shor in the Experimental Regime
@@
A-G Effect of Decoherence on the QPE Output Distribution

### v2

@@ 6 Trends in Quantum Computing Developments
Table 4: Company roadmaps with numbers of physical or logical qubits.
@@
Trapped-ion providers such as IonQ and Quantinuum focus on saller but higher-fidelity systems. In 2024, Quantinuum introduced a 56-qubit system that uses ion shuttling across a chip to perform operations pushing system level benchmarks such as quantum volume (QV) [34, 35] to new record values. In 2025, the company released a 98 qubit next-generation system, which reached a point where the QV benchmark becomes too expensive in terms of classical computation to be carried out. IonQ targets 64 qubits by 2026 and 1,024 qubits by 2028 by connecting separate chips via photonic links for scalability. Their recent acquisition of smaller startup Oxford Ionics makes them the record holder in terms of fidelities with a two-qubit gate above $99.99\%$ fidelity [36] and SPAM fidelities as high as $99.9993\%$ [37], albeit this is for an architecture involving microwave control different to their original laser-based one.
Photonic-qubit companies, notably PsiQuantum and Xanadu, pursue architectures that exploit integrated optics and telecom technologies. PsiQuantum has proposed a path to a one-million-qubit fault-tolerant computer by 2029, based on a fusion-based architecture [38] that combines small photonic resource states in a probabilistic manner. However, major technical milestones, including the realization of large-scale, on-chip entanglement with low loss, are still outstanding. Xanadu is taking a different approach by using squeezed states of light, which are more resistant to signal loss. In 2022, the company showcased a prototype device that achieved a milestone similar to Google’s breakthrough in quantum computational advantage [39]. In addition to hardware, Xanadu develops the open-source PennyLane platform to support hybrid quantum-classical algorithm design.
@@
For context, earlier resource estimates suggest that factoring a 2048-bit RSA modulus using a relatively direct implementation of Shor’s algorithm would require on the order of 8.4 million physical qubits [40], assuming that they can be operated with sufficiently low error rates and long coherence times. More recent analyses, which incorporate algorithmic improvements and more efficient QEC pipelines, reduce this requirement but still call for very large, high-quality devices [21, 22, 23].
@@
The best reported error rates (all in trapped ions) for single qubit operations are currently at the level of $10^{-7}$ [41], with the best two-qubit gates having reached the $10^{-5}$ range [36]. These numbers, however, are “hero values”. In many multi-qubit systems, effective error rates are much closer to $10^{-4}$ for single-qubit and $10^{-2}$ for two-qubit gates once full-system effects are included. In contrast, modern digital transistors exhibit error rates as low as $10^{-23}$ [42], highlighting the vast gap between classical and quantum reliability.
Large-scale, fault-tolerant quantum computation therefore hinges on effective QEC. In recent years, experimental demonstrations have shown that QEC can, in principle, suppress logical error rates below those of the underlying physical qubits. For instance, a demonstration in 2025 [43] reported an experiment in which an error metric is reduced from $10^{-2}$ to $10^{-7}$ using 72 physical qubits. However, that demonstration relied on post-processing rather than real-time feedback and did not correct all error channels. Real-time, active QEC remains experimentally demanding and has so far only been achieved in limited settings, e.g. in a trapped-ion experiment [44].
@@ A.2 Per-Run Success Probability
A.3 Importance of the Quantum Signal
@@
A.4 Certifying Substantial Quantum Signal in Practice
@@
A.5 Shor in the Ideal Regime
@@
A.6 Shor in the Experimental Regime
@@
A.7 Effect of Decoherence on the QPE Output Distribution